# Compositional Reasoning with Transformers, RNNs, and Chain of Thought

### v1

@@ D.2 Memory-rank for non-binary trees
In this sub-section we explain how to extend some of the results from Section 5 to non-binary trees. We first define memory-rank for general trees:
@@
It is straightforward to see that this definition generalizes Definition 5.3. Also, Algorithm 1 readily generalizes to Definition D.1, since it has no dependence on the degree of each node.
Extending Theorem 5.4 to general trees with maximal degree of $k$ can be done in a straightforward way by having an RNN with hidden dimension of $O(dk\log(n))$ and $O(k)$ layers. The idea is to add for each node an additional coordinate representing its degree. Now, in the proof of Theorem 5.4, when the last two nodes are extracted from the stack (layer $3$ in the construction), we split this into $k$ layers. Each layer checks the degree of the parent node from $2$ until $k$ and for degree $i$ extracts the last $i$ nodes from the stack. This can be implemented by a ReLU network of depth $O(k)$. Finally, the node with the highest inner product to the parent is inserted back to the stack with an updated depth (similarly to layers $4$ and $5$ from the proof of Theorem 5.4).
We believe it is also possible to extend this construction to having depth $O(1)$ and hidden dimension $O(d\log(n))$ for general trees by changing the sorting algorithm. This can be done by splitting the argmax operation over $k$ inputs into $k-1$ argmax operations, each over $2$ inputs, where these operations are nested. We leave this construction for future works.
Appendix E Proofs from Section 6
The main idea of the construction is to iteratively use the self-attention layer constructed in Theorem 4.1, while wrapping it with a construction that allows outputting CoT tokens in the right order.
By Lemma A.1 there exist vectors $\mathbf{z}_{1},\dots,\mathbf{z}_{n}\in\{\pm 1\}^{4\log(n)}$ such that $|\left\langle\mathbf{z}_{i},\mathbf{z}_{j}\right\rangle|\leq 3\log(n)$ for any $i\neq j$. We will use these vectors twice, and for two different use cases. The first one is similar to the construction of Theorem 4.1. Namely, for each node $v\in V$ that is not a leaf, we set one of the vectors $\mathbf{z}_{i}$ to be corresponded with this node and denote it as $\mathbf{z}_{v}$.
We also assume that the tree has a reverse BFS ordering. This means that we do a standard BFS ordering of the tree, and order of node $i$ in the reverse order is $n-i+1$. For a node $v$, let $I(v)$ be be its place in the reverse BFS ordering. We also correspond one of the vectors defined above to each place in this order we denote this vector as $\mathbf{w}_{I(v)}$. For the root $v_{r}$ we have $I(v_{r})=n$, and define $\mathbf{w}_{I(v_{r})+1}=\mathbf{w}_{1}$. We use a different notation for the BFS ordering to not confuse them with the other set of positional encoding, although this is the same set of vectors. It is also possible to find another set of vectors with the same property, although it won’t matter for the construction. The embedding for a leaf node $u$ and a non-leaf node $v$ are defined as:
@@ Appendix E Proofs from Section 6
As in Theorem 4.1, all the coordinates of the positional encoding are in ±$\pm\{1\}$, and the matrices that operate over these coordinates will have all their entries in $\{-1,0,1\}$. We will first give an intuitive explanation of the embedding and how the transformer will work. The embedding first contains the vector $\mathbf{x}$ corresponding to the node. It also contains its own positional encoding $\mathbf{z}_{v}$ if its a non-leaf, and the positional encoding of its parent. Next, it contains the vector representing its place in the reverse BFS ordering, as well as the subsequent vector in the order. The last coordinate is a flag for whether a node is a leaf.
The transformer will contains $2$ self-attention layers, and an MLP with $O(1)$ layers after each self-attention layer. The first self-attention layer and the MLP afterwards is similar to the construction in Theorem 4.1. Their goal is to solve the sub-questions in the CRQ. This layer will use the $\mathbf{z}$’s positional encoding vectors. The second self-attention layer will make sure that the last CoT token that was created will attend to the next token corresponding to the node after it in the reverse BFS ordering. This way, each CoT token will provide an answer to one sub-question of the CRQ, and using the reverse BFS order we make sure that any subsequent questions will have the questions of their children already solved.
@@
We now turn to prove the correctness of the construction. The second self-attention layer forces each token $v$ to attend to the token $u$ with $I(u)=I(v)+1$, namely the consecutive token in the reverse BFS order. The $V$ matrix of the second self-attention layer transform the embedding of this token to be similar to that of an embedding of a node. Suppose we constructed $k-1$ CoT tokens already, and for a node $v$, $I(v)=k$. By the ordering, this means that all the descendants of $v$ have being processed. Also, if $v$ have descendants that are non-leaves, then the first layer correctly solved the corresponding sub-question to them, and they were outputted as CoT tokens with a similar embedding to a leaf node. Hence, in the current iteration, the token corresponding to $v$ will necessarily attend the token corresponding to its child that solves the sub-question. In addition, the last token CoT token that was created will necessarily attend to the token corresponding to $v$ by the construction of the second self-attention layer. This means that the new CoT token will be the answer to the sub-question corresponding to $v$, with and embedding similar to a leaf. This is true for any node $v$, hence the last CoT token will contain the answer to the root, which is the answer to the CRQ.
@@ Appendix F General CRQs
In this appendix we present a generalization of CRQs by allowing different operators. For simplicity we only consider binary trees. Given a CRQ over a non-binary tree we can always define an equivalent CRQ over a binary tree, where the number of nodes increase by a factor of at most $2$, see Appendix D. We begin with a formal definition:
@@
A General Compositional Reasoning Question (GCRQ) is a rooted binary tree $T=(V,E)$ with root $v_{r}\in V$, and a set of operators $f_{1},\dots,f_{k}:\mathbb{R}^{3d}\rightarrow\mathbb{R}^{d}$. Each node $v\in V$ is labeled by a vector $\mathbf{x}_{v}\in\Gamma^{d}$, and non-leaf nodes are also labeled by a number $i_{v}\in[k]$. Here $\Gamma\subset\mathbb{R}$ is some finite vocabulary of constant size and $d\in\mathbb{N}$.
@@
For a concrete example, we show how to implement arithmetic problems using GCRQs. Let $\Gamma=\{0,\pm 1,\dots,\pm A\}$ for some constant integer $A$. Also assume that $d=1$. We define the operators: $f_{+},f_{-},f_{\times},f_{\div}$ as:
@@
Note that each such operator can be represented by a ReLU network. For $f_{+}$ and $f_{-}$ this is clear, since these are linear operators. For $f_{\times}$ we can use Lemma C.3 to approximate $(x,y)\mapsto x\cdot y$ up to an accuracy of $\frac{1}{4}$ with a $2$-layer network. Since the alphabet is finite, We can threshold over integers using another ReLU layer to get an exact representation of multiplication over a finite set of integers. A similar construction can be used for $f_{\div}$.
@@
We now explain how to change our constructions to work with GCRQs. Note that the only place in the proofs of Theorem 4.1 and Theorem 6.1 we use the fact that the operator is argmax, is where we use a single self-attention layer to calculate inner products. Instead, we can use such a layer so that the parent node will attend to both its children with the output being a concatenation of their vectors. In more details, assume that each two siblings $u_{1}$ and $u_{2}$ in the tree are labeled by vectors $\begin{pmatrix}\mathbf{x}_{u_{1}}\\ \boldsymbol{0}_{d}\end{pmatrix}$ and $\begin{pmatrix}\boldsymbol{0}_{d}\\ \mathbf{x}_{u_{2}}\end{pmatrix}$. Then, their parent can attend to both of them with the same weight using the positional embeddings, in the same way as done in the proofs of the mentioned theorems. We also construct the MLP so that its input will include the reference to the operator $i_{v}$. The MLP will simulate an ”if“ statement over the $k$ operators and apply the relevant one over the vectors $\mathbf{x}_{v},\mathbf{x}_{u_{1}},\mathbf{x}_{u_{2}}$. This ”if“ statement can be simulated by a $k$-layer ReLU network (we assume the $k$ is independent of $n$).
@@
A similar construction will work for the proof of Theorem 5.4, where here the MLP will output the last two inputs from the stack and apply the operator $i_{v}$. Again, it will increase the depth of the MLP by $k$. We leave exact constructions for future work.

### v2

@@ D.2 Memory-rank for non-binary trees
In this sub-section we explain how to extend some of the results from Section˜5 to non-binary trees. We first define memory-rank for general trees:
@@
It is straightforward to see that this definition generalizes Definition˜5.3. Also, Algorithm˜1 readily generalizes to Definition˜D.1, since it has no dependence on the degree of each node.
Extending Theorem˜5.4 to general trees with maximal degree of $k$ can be done in a straightforward way by having an RNN with hidden dimension of $O(dk\log(n))$ and $O(k)$ layers. The idea is to add for each node an additional coordinate representing its degree. Now, in the proof of Theorem˜5.4, when the last two nodes are extracted from the stack (layer $3$ in the construction), we split this into $k$ layers. Each layer checks the degree of the parent node from $2$ until $k$ and for degree $i$ extracts the last $i$ nodes from the stack. This can be implemented by a ReLU network of depth $O(k)$. Finally, the node with the highest inner product to the parent is inserted back to the stack with an updated depth (similarly to layers $4$ and $5$ from the proof of Theorem˜5.4).
We believe it is also possible to extend this construction to having depth $O(1)$ and hidden dimension $O(d\log(n))$ for general trees by changing the sorting algorithm. This can be done by splitting the argmax operation over $k$ inputs into k1k-1 argmax operations, each over $2$ inputs, where these operations are nested. We leave this construction for future works.
Appendix E Proofs from Section˜6
The main idea of the construction is to iteratively use the self-attention layer constructed in Theorem˜4.1, while wrapping it with a construction that allows outputting CoT tokens in the right order.
By Lemma˜A.1 there exist vectors $\mathbf{z}_{1},\dots,\mathbf{z}_{n}\in\{\pm 1\}^{4\log(n)}$ such that $|\left\langle\mathbf{z}_{i},\mathbf{z}_{j}\right\rangle|\leq 3\log(n)$ for any $i\neq j$. We will use these vectors twice, and for two different use cases. The first one is similar to the construction of Theorem˜4.1. Namely, for each node $v\in V$ that is not a leaf, we set one of the vectors $\mathbf{z}_{i}$ to be corresponded with this node and denote it as $\mathbf{z}_{v}$.
We also assume that the tree has a reverse BFS ordering. This means that we do a standard BFS ordering of the tree, and order of node $i$ in the reverse order is ni+1n-i+1. For a node $v$, let $I(v)$ be be its place in the reverse BFS ordering. We also correspond one of the vectors defined above to each place in this order we denote this vector as $\mathbf{w}_{I(v)}$. For the root $v_{r}$ we have $I(v_{r})=n$, and define $\mathbf{w}_{I(v_{r})+1}=\mathbf{w}_{1}$. We use a different notation for the BFS ordering to not confuse them with the other set of positional encoding, although this is the same set of vectors. It is also possible to find another set of vectors with the same property, although it won’t matter for the construction. The embedding for a leaf node $u$ and a non-leaf node $v$ are defined as:
@@ Appendix E Proofs from Section˜6
As in Theorem˜4.1, all the coordinates of the positional encoding are in $\{\pm 1\}$, and the matrices that operate over these coordinates will have all their entries in $\{-1,0,1\}$. We will first give an intuitive explanation of the embedding and how the transformer will work. The embedding first contains the vector $\mathbf{x}$ corresponding to the node. It also contains its own positional encoding $\mathbf{z}_{v}$ if its a non-leaf, and the positional encoding of its parent. Next, it contains the vector representing its place in the reverse BFS ordering, as well as the subsequent vector in the order. The last coordinate is a flag for whether a node is a leaf.
The transformer will contain $2$ self-attention layers, and an MLP with $O(1)$ layers after each self-attention layer. The first self-attention layer and the MLP afterwards is similar to the construction in Theorem˜4.1. Their goal is to solve the sub-questions in the CRQ. This layer will use the $\mathbf{z}$’s positional encoding vectors. The second self-attention layer will make sure that the last CoT token that was created will attend to the next token corresponding to the node after it in the reverse BFS ordering. This way, each CoT token will provide an answer to one sub-question of the CRQ, and using the reverse BFS order we make sure that any subsequent questions will have the questions of their children already solved.
@@
We now turn to prove the correctness of the construction. The second self-attention layer forces each token $v$ to attend to the token $u$ with $I(u)=I(v)+1$, namely the consecutive token in the reverse BFS order. The $V$ matrix of the second self-attention layer transform the embedding of this token to be similar to that of an embedding of a node. Suppose we constructed k1k-1 CoT tokens already, and for a node $v$, $I(v)=k$. By the ordering, this means that all the descendants of $v$ have being processed. Also, if $v$ have descendants that are non-leaves, then the first layer correctly solved the corresponding sub-question to them, and they were outputted as CoT tokens with a similar embedding to a leaf node. Hence, in the current iteration, the token corresponding to $v$ will necessarily attend the token corresponding to its child that solves the sub-question. In addition, the last token CoT token that was created will necessarily attend to the token corresponding to $v$ by the construction of the second self-attention layer. This means that the new CoT token will be the answer to the sub-question corresponding to $v$, with and embedding similar to a leaf. This is true for any node $v$, hence the last CoT token will contain the answer to the root, which is the answer to the CRQ.
@@ Appendix F General CRQs
In our definition of CRQs (Definition˜3.1) each sub-question is defined as the argmax over inner products. There are two main reasons for this choice: (1) The argmax function aligns well with the attention layers in transformers, allowing for simple constructions, and (2) The argmax function is flexible enough to capture many natural problems as special cases, such as evaluating Boolean expressions. However, by allowing for other operations besides argmax, CRQs can be generalized to include many more tasks. We define the class of generalized CRQ problems formally below. For simplicity we only consider binary trees. (Given a CRQ over a non-binary tree, we can always convert it into an equivalent CRQ over a binary tree with at most twice as many nodes; see Appendix˜D.)
@@
A General Compositional Reasoning Question (GCRQ) is a rooted binary tree $T=(V,E)$ with root $v_{r}\in V$, and a set of operators $f_{1},\dots,f_{k}:\mathbb{R}^{3d}\rightarrow\mathbb{R}^{d}$. Each node $v\in V$ is labeled by a vector $\mathbf{x}_{v}\in\Gamma^{d}$, and non-leaf nodes are also labeled by an operator’s index $i_{v}\in[k]$. Here $\Gamma\subset\mathbb{R}$ is some finite vocabulary of constant size and $d\in\mathbb{N}$.
@@
As a concrete example, we can represent modular arithmetic problems as a special case of GCRQs. Let $\Gamma=\{0,1,\dots,p-1\}$ for some prime $p$. Let the dimension of the embedding vectors be $d=1$. We define the operators: $f_{+},f_{-},f_{\times},f_{\div}$ as:
@@
(In this construction, the non-leaf labels $\mathbf{x}_{v}$ are not used.) Clearly, any arithmetic expression can be converted into a GCRQ whose tree structure matches the order of operations in the expression. Furthermore, each of the four arithmetic operators can be represented by a ReLU network. For instance, Feng et al. [2024, Lemma C.5, Theorem D.1] implements them using a lookup table, since the inputs come from a finite field.
@@
All of our constructions can be adapted to work for GCRQs. Consider $k$ operators $f_{1},\dots,f_{k}:\mathbb{R}^{3d}\rightarrow\mathbb{R}^{d}$, where each operator can be implemented by a ReLU network. Note that the only place in the proofs of Theorem˜4.1 and Theorem˜6.1 we use the fact that the operator is argmax is where we use a single self-attention layer to calculate inner products. Instead, we can modify this attention layer so that each parent node attends to both its children, with the output being a concatenation of their vectors. In more detail, assume that each two siblings $u_{1}$ and $u_{2}$ in the tree are labeled by vectors $\begin{pmatrix}\mathbf{x}_{u_{1}}\\ \boldsymbol{0}_{d}\end{pmatrix}$ and $\begin{pmatrix}\boldsymbol{0}_{d}\\ \mathbf{x}_{u_{2}}\end{pmatrix}$. Then using the positional embeddings, we can force their parent to attend to both of them with the same attention score just as in the proofs of Theorem˜4.1 and Theorem˜6.1. We also construct the MLP so that its input will include the index of the operator, $i_{v}$. The MLP will simulate an ”if“ statement over the $k$ operators and apply the relevant one over the vectors $\mathbf{x}_{v},\mathbf{x}_{u_{1}},\mathbf{x}_{u_{2}}$. This ”if“ statement can be simulated by a $k$-layer ReLU network. We need to assume that $k$ is independent of $n$, which is natural since the number of possible operators (e.g. $4$ in arithmetic problems) is independent of the input length.
@@
A similar construction will work for the proof of Theorem˜5.4, where here the MLP will output the last two inputs from the stack and apply the operator $i_{v}$. Again, it will increase the depth of the MLP by $k$. We leave exact constructions for future work.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our abstract and introduction highlight the main results of each section: the definition of CRQs, the lower bounds (Theorem˜4.3, Theorem˜5.5), and the upper bounds (Theorem˜4.1, Theorem˜5.4, Theorem˜6.1).
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Each of our theorems clearly states all assumptions. Section˜7 discusses the questions raised by our work that remain open.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Formal theorem statements include all assumptions. Proofs appear in Appendices˜A, B, C and E.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [N/A]
  19. Justification: This paper does not include experiments.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [N/A]
  24. Justification: This paper does not include experiments.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [N/A]
  29. Justification: This paper does not include experiments.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [N/A]
  34. Justification: This paper does not include experiments.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [N/A]
  39. Justification: This paper does not include experiments.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: No human subjects or data was used. No clear potential exists for harmful consequences.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [No]
  49. Justification: This paper provides theoretical comparison of several widely-used LLM architectures. The goal is to achieve better scientific understanding of these methods. While LLMs have broad societal impacts, the results of this paper do not directly relate to this impact.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: No data or models are released.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [N/A]
  59. Justification: No such existing assets are used.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: No new assets are released.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: No research with human subjects was performed.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: No research with human subjects was performed.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLMs were not used in any meaningful way.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.